Suppose x is a normally distributed random variable with μ = 50 and Ϭ = 3. Find a value of the random variable, call it x0, such that a. P(x ≤ x0) = 0.8413 b. P(x > x0) = 0.25 c. P(x > x0) = 0.95 d. P(41 ≤ x < x0) = 0.8630 e. 10% of the values of x are less than x0. f. 1% of the values of x are greater than x0.
1
Expert's answer
2020-07-29T14:44:00-0400
"x\sim N(\mu, \sigma^2)"

Given "\mu=50, \sigma=3"

a.

"P(x\leq x_0)=0.8413"

"P(x\leq x_0)=P(z\leq z_0 )=0.8413"

"z_0=0.999815={x_0-\mu \over \sigma}"

"x_0=3(0.999815)+50\approx53"

b.

"P(x>x_0)=0.25"

"P(x\leq x_0)=1-P(x>x_0)=1-0.25=0.75"

"P(x\leq x_0)=P(z\leq z_0 )=0.75"

"z_0=0.674490={x_0-\mu \over \sigma}"

"x_0=3(0.674490)+50\approx52"

c.


"P(x>x_0)=0.95"

"P(x\leq x_0)=1-P(x>x_0)=1-0.95=0.05"

"P(x\leq x_0)=P(z\leq z_0 )=0.05"

"z_0=-1.645={x_0-\mu \over \sigma}"

"x_0=3(-1.645)+50\approx45"

d.


"P(41\leq x< x_0)=P(x<x_0)-P(x\leq 41)="

"=P(z<z_0)-P(z\leq{41-50 \over 3})="

"=P(z<z_0)-P(z\leq-3)="

"=P(z<z_0)-0.00135=0.8630"

"P(z<z_0)=0.86435"

"z_0=1.1={x_0-\mu \over \sigma}"

"x_0=3(1.1)+50\approx53.3"

e.


"P(x< x_0)=0.1"

"P(x< x_0)=P(z<z_0)=0.1"

"z_0=-1.28155={x_0-\mu \over \sigma}"

"x_0=3(-1.28155)+50\approx46.155"

f.


"P(x>x_0)=0.01"

"P(x\leq x_0)=1-P(x>x_0)=1-0.01=0.99"

"P(x\leq x_0)=P(z\leq z_0 )=0.99"

"z_0=2.326={x_0-\mu \over \sigma}"

"x_0=3(2.326)+50\approx56.978"